Suppose you select 2 letters at random from the word "pacific". find each probability
P(2 vowels)
P(2 consonants)
P(1 vowel, 1 consonant)

Answer by stanbon(75887)

Suppose you select 2 letters at random from the word "pacific".
vowels: a,i,i
consonants: p,c,f,c
-----------
find each probability
P(2 vowels) = 3C2/7C2 = 3/21 = 1/7
--------------
P(2 consonants) = 4C2/7C2 = 6/21 = 2/7
--------------
P(1 vowel, 1 consonant) = 1 - [1/7 + 2/7] = 4/7
=====================================================
Cheers,
Stan H.